what is is the value of x and y in the simultaneous equation 2x+27=4
3x+3y=6

Answer by checkley77(12844)

2x+2(y)=4 multiply by 3
3x+3y=6 multiply by -2 & add.
6x+6y=12
-6x-6y=-12
---------------
0x+0x=0 thus no unique solution to these 2 equations.
They represent the same line with infinite solutions.